# Feynman integrals at large loop order and the $\log$-$\Gamma$ distribution

### v2

@@ 1 Introduction
Feynman integrals are the building blocks of perturbative quantum, statistical, and classical field theory expansions. Each Feynman integral contributes to a specific perturbative order. In most cases, this order equals the graph’s loop number. A perturbative expansion is a formal power series $\mathcal{A}(\hbar)=\sum_{L\geq 0}A_{L}\hbar^{L}$, with the $L$-th coefficient given by a sum of Feynman integrals:
@@
where $\hbar$ is the perturbative expansion parameter, we sum over all Feynman graphs $G$ of specific shape and loop order $L(G)$, $I_{G}$ is the Feynman integral corresponding to the graph, and $|\operatorname{Aut}(G)|$ denotes the symmetry factor of the graph (i.e. the order of its automorphism group). The precise shape of the graphs and the associated integral depend on the specific underlying theory.
@@
Ultimately, this question is motivated by the perturbation theory at large order program [12, 13] and its recent incarnation, the resurgence program (see, e.g., [14]). The progress from both these programs suggests that the large-$L$ behavior of the coefficients $A_{L}$ encodes much (and perhaps all) nonperturbative information of the function $\mathcal{A}(\hbar)$. Further, an ongoing research program aims to replace sums as (1) with integrals over a single, continuous object [15, 16, 17, 18]. Our results provide concrete information on such an object’s expected asymptotic $L\rightarrow\infty$ shape.
This article will focus on a specific model: $\phi^{4}$ quantum field theory in four-dimensional spacetime. Within this model, we focus on a particular observable: the primitive contribution to the beta function. The $\phi^{4}$ beta function is known exactly up to loop order $7$ in the minimal subtraction (MS) scheme [19, 20, 21, 4]. The primitive contribution to this beta function is obtained by summing over all period Feynman integrals, given by the $1/\varepsilon$ residues of specific Feynman integrals (see §2 for a precise definition of period Feynman integrals). It is conjectured that asymptotically, at large loop order, the primitive contribution gives the dominant contribution to the beta function of $\phi^{4}$ theory in the MS scheme [22] (see also [23, 24, 25]).
In §2, we use an empirical, numerical approach to study the terms in the sum (1) when $L$ is large. At sufficiently large loop order, exact computation methods for Feynman integrals will inevitably fail. Here, we use the tropical sampling approach introduced by the first author in [26] to evaluate many Feynman integrals with up to 17 loops (see also [27]). The tropical sampling method draws from previous ideas of sector decomposition [28, 29] and the Hepp bound [30]. Balduf and Balduf–Shaban recently performed similar large-scale computations of Feynman integrals using the tropical approach [31, 32]. In [31], the primitive contribution to the $\phi^{4}$ theory beta function was estimated up to 18 loops. We confirm these computations up to 17 loops.
@@
The histogram shows our measured distribution of $\phi^{4}$ period Feynman integrals at 17 loops. We obtained this distribution by evaluating $44027$ Feynman integrals up to $10^{-3}$ relative accuracy. Histograms similar to Figure 1 already appeared in [33] (see also [34]), which studied the distribution of Feynman integrals with a related underlying motivation. Here, thanks to more advanced tools, we have the advantage of being able to probe a much higher order in perturbation theory. Below, Figure 2 shows similar histograms for lower loop orders and illustrates the rapid convergence to the limiting distribution.
The red curve depicts the density
@@
where $\mu(\mathcal{P})$ is the density of period Feynman integrals with value $\mathcal{P}$, and the parameters $\alpha=6.41(5)$, $\lambda=3.87(3)$, and ${{\mathcal{P}_{0}}=5.14(7)\cdot 10^{4}}$ are fitted and specific to $L=17$ (see §3.1 for details on the fitting method). The density (2) is called a $\log$-$\Gamma$ distribution because, as a function of $\log\mathcal{P}$, it is the shifted integrand of Euler’s representation for the $\Gamma$ function. This motivates
@@
The distribution of period Feynman integrals in $\phi^{4}$ theory at $L$ loops weakly converges to a $\log$-$\Gamma$ distribution when $L\rightarrow\infty$.
@@
In general, the coefficients $A_{L}$ are expected to grow factorially [35] and, hence, the perturbative series, such as $\mathcal{A}(\hbar)=\sum_{L\geq 0}A_{L}\hbar^{L}$, to be divergent if $\hbar\neq 0$. To still utilize the information of the coefficients $A_{L}$ and to rigorously compute a quantity such as $\mathcal{A}(\hbar)$ for nonzero $\hbar$, resummation techniques are needed. As input data, these techniques need information about the $L\rightarrow\infty$ asymptotic behavior of $A_{L}$. For this reason, this asymptotic behavior is of high conceptual and practical value for the perturbative quantum field theory framework [36, 37, 38, 39, 40]. One concrete application is the prediction of critical exponents of various interesting universality classes via the Wilson–Fischer approach [2] (see also [41] for a recent discussion of the relevance of large-order contributions in this domain). In §3, we tabulate our 17 loop order results for the value of the primitive contribution to the $\phi^{4}$ beta function. We then extrapolate these values to the limit $L\rightarrow\infty$ and discuss the relation to classic conjectures on the asymptotic growth rate of the $\phi^{4}$ beta function in the MS scheme (e.g. to [42, 22, 24, 25]). We conclude in §4.
@@ 2.1 The primitive contribution to the $\phi^{4}$ beta function
In four dimensions, the superficial degree of divergence of a scalar Feynman graph is given by $\omega(G)=|E_{G}|-2L(G)$, where $|E_{G}|$ is the number of edges of the graph and $L(G)$ its loop number [43]. A graph $G$ is primitive divergent in $\phi^{4}$ theory if it is 1PI, has precisely four external legs, and $\omega(\gamma)>0$ for each proper subgraph $\gamma\subsetneq G$, while $\omega(G)=0$. The last conditions ensure the graph has an overall logarithmic divergence and no subdivergences in four-dimensional spacetime. As usual, we will consider the legs of the graphs fixed or equivalently distinguishable.
@@
Our main example is the primitive contribution to the $\phi^{4}$ beta function (see [23, Appendix B], whose notation we follow, for details). This contribution is renormalization group independent and therefore appears in every renormalization scheme. We may express this contribution in terms of period Feynman integrals,111 The convential shift by one in $\beta^{\mathrm{prim}}_{L+1}$ is due to the fact that $L$-loop vertex diagrams contribute to the $g^{L+1}$ coefficient of the beta function in the coupling $g$.
@@ 2.2 A probabilistic approach to the sum over graphs
As the legs are distinguishable (which can be realized on the computer by numbering them, for example), the graph returned by the algorithm will have four distinguishable legs. It follows from a simple combinatorial argument and the orbit stabilizer theorem that the algorithm above randomly produces graphs $G$ with probability $p(G)$ from (6). Combining (5) and (6), gives
@@ 3.1 Histograms of Feynman integrals at large loop order
where we sum over all period samples $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}$ at a specific loop order. This likelihood function is readily derived by multiplying $N$ copies of the function that multiplies the measure in (2) and taking the logarithm. As the number of samples $N$ is large, we can estimate the uncertainties of these parameters by approximating the prior distribution using a Gaussian. The resulting parameters with uncertainties are listed in Table 1. The uncertainties of the fit parameters were extremely large for $L\in\{8,9\}$. So, we discarded these fits. The fitted distributions are depicted as red lines in Figure 1 and Figure 2.
@@
We checked Conjecture 1 quantitatively using Pearson’s $\chi^{2}$ test: Let $O_{i}$ be the number of evaluated period Feynman integrals that fall into the $i$-th percentile of the distribution (2) with the fitted maximum likelihood parameters at the respective loop order. The expectation value of the random variable $O_{i}$ is $N/100$. So, under the hypothesis that our data follows (2), the quantity $\chi^{2}=\frac{100}{N}\sum_{i}(O_{i}-\frac{N}{100})^{2}$ is expected to follow a $\chi^{2}$-distribution with mean $100-3=97$, as three parameters are fitted. Table 1 shows that the ratio $\chi^{2}/97$ approaches $1$ with increasing loop order, consistent with Conjecture 1. Figure 2 illustrates how the distribution is approached with increasing loop number, providing further evidence for Conjecture 1.
@@
$L$ $\alpha$ $\lambda$ $\mathcal{P}_{0}$ $\frac{\chi^{2}}{97}$ $\langle\mathcal{P}(G)\rangle_{L}$ $\beta_{L+1}^{\rm prim}$    $N$
$8$ $7.449(7)\cdot 10^{2}$ $6.064(5)\cdot 10^{6}$ $99900$
$9$ $1.644(2)\cdot 10^{3}$ $1.046(1)\cdot 10^{8}$ $99800$
$10$ $19.72(84)$ $10.6(2)$ $4.97(19)\cdot 10^{2}$ $273.7$ $3.504(5)\cdot 10^{3}$ $1.890(3)\cdot 10^{9}$ $99900$
$11$ $10.69(27)$ $6.9(1)$ $1.36(3)\cdot 10^{3}$ $67.7$ $7.237(12)\cdot 10^{3}$ $3.558(6)\cdot 10^{10}$ $99700$
$12$ $8.36(17)$ $5.53(6)$ $2.78(4)\cdot 10^{3}$ $19.3$ $1.460(3)\cdot 10^{4}$ $6.998(14)\cdot 10^{11}$ $100000$
$13$ $7.38(13)$ $4.85(5)$ $5.28(6)\cdot 10^{3}$ $7.3$ $2.869(6)\cdot 10^{4}$ $1.429(3)\cdot 10^{13}$ $98712$
$14$ $6.88(11)$ $4.42(4)$ $9.58(10)\cdot 10^{3}$ $4.6$ $5.531(14)\cdot 10^{4}$ $3.036(7)\cdot 10^{14}$ $100000$
$15$ $6.83(10)$ $4.24(4)$ $1.67(2)\cdot 10^{4}$ $2.5$ $1.039(3)\cdot 10^{5}$ $6.654(18)\cdot 10^{15}$ $100000$
$16$ $6.688(96)$ $4.08(4)$ $2.92(3)\cdot 10^{4}$ $2.2$ $1.910(6)\cdot 10^{5}$ $1.504(4)\cdot 10^{17}$ $99650$
$17$ $6.41(12)$ $3.87(5)$ $5.14(7)\cdot 10^{4}$ $1.2$ $3.489(17)\cdot 10^{5}$ $3.556(17)\cdot 10^{18}$ $44027$
@@
Table 1: Fit parameters $\alpha,\lambda,$ and $\mathcal{P}_{0}$ for the period Feynman integral distribution in $\phi^{4}$ theory, the normalized $\chi^{2}$ value resulting from Pearson’s $\chi^{2}$ test, the estimated average value of the period Feynman integral, the estimated value of the primitive beta function coefficient, and the number $N$ of Feynman integrals evaluated at each loop order.

### v3

@@ 1 Introduction
Feynman integrals are the building blocks of perturbative quantum, statistical, and classical field theory expansions. Each Feynman integral corresponds to a Feynman graph and contributes to a specific perturbative order. In most cases, this order equals the graph’s loop number. A perturbative expansion is a formal power series $\mathcal{A}(\hbar)=\sum_{L\geq 0}A_{L}\hbar^{L}$, with the $L$-th coefficient given by a sum of Feynman graphs:
@@
where $\hbar$ is the perturbative expansion parameter, we sum over all Feynman graphs $G$ of specific shape and loop order $L(G)$, $I_{G}$ is the Feynman integral corresponding to the graph, and $|\operatorname{Aut}(G)|$ denotes the graph’s symmetry factor (i.e. the order of its automorphism group). The precise shape of the graphs and the associated integral depend on the specific underlying theory.
@@
Ultimately, this question is motivated by the perturbation theory at large order program [12, 13] and its recent incarnation, the resurgence program (see, e.g., [14]). The progress from both these programs suggests that the large-$L$ behavior of the coefficients $A_{L}$ encodes much (and perhaps all) nonperturbative information of the function $\mathcal{A}(\hbar)$. Further, an ongoing research program aims to replace sums as (1) with integrals over a single, continuous object [15, 16, 17, 18]. Our results in Sec. 3.1 provide concrete information on such an object’s expected asymptotic $L\rightarrow\infty$ shape.
This article will focus on a specific model: $\phi^{4}$ quantum field theory in $D=4-2\varepsilon$ dimensions. Within this model, we focus on a particular subset of terms contributing to the beta function: the primitive contribution. The $\phi^{4}$ beta function is known exactly up to loop order $7$ in the minimal subtraction (MS) scheme [19, 20, 21, 4]. The primitive contribution to this beta function is obtained by summing over all period Feynman integrals, given by the $1/\varepsilon$ residues of specific Feynman integrals that are free of subdivergences (see §2 for a precise definition of period Feynman integrals). It is conjectured that asymptotically, at large loop order, the primitive contribution gives the full beta function of $\phi^{4}$ theory in the MS scheme [22] and other contributions are subdominant in that limit (see also [23, 24, 25]).
In §2, we use an empirical, numerical approach to study the terms in the sum (1) when $L$ is large. At sufficiently large loop order, exact computation methods for general Feynman integrals will inevitably fail. Here, we use the tropical sampling approach introduced by the first author in [26] to evaluate many Feynman integrals with up to 17 loops (see also [27]). The tropical sampling method draws from previous ideas of sector decomposition [28, 29], the Hepp bound [30], and the algebraic geometry of Feynman integrals [9, 31], which it leverages with properties of generalized permutahedra [32]. Balduf and Balduf–Shaban recently performed similar large-scale computations of Feynman integrals using the tropical approach [33, 34]. Balduf estimated the primitive contribution to the $\phi^{4}$-theory beta function up to 18 loops [33]. We confirm these computations up to 17 loops.
@@
The histogram shows our measured distribution of $\phi^{4}$ period Feynman integrals at 17 loops. We obtained this distribution by evaluating $44027$ Feynman integrals up to $10^{-3}$ relative accuracy. Histograms similar to Figure 1 already appeared in [35] (see also [36]), which studied the distribution of Feynman integrals with a related underlying motivation. Here, thanks to more advanced tools, we have the advantage of being able to probe a much higher order in perturbation theory. Below, Figure 2 shows similar histograms for lower loop orders and illustrates the rapid convergence to the limiting distribution. The red curve depicts the density
@@
where $\mu(\mathcal{P})$ is the density of period Feynman integrals with value $\mathcal{P}$, and the parameters $\alpha=6.41(5)$, $\lambda=3.87(3)$, and ${\mathcal{P}_{0}}=5.14(7)\cdot 10^{4}$} are fitted and specific to $L=17$ (see §3.1 for details on the fitting method). The density (2) is called a $\log$-$\Gamma$ distribution (see, e.g., [37, Example 18.4]) because, as a function of $\log\mathcal{P}$, it is the shifted integrand of Euler’s representation for the $\Gamma$ function, $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\text{d}t.$ This motivates
@@
The distribution of period Feynman integrals in $\phi^{4}$ theory at $L$ loops sampled with a probability inversely proportional to the symmetry factor of the associated Feynman graph weakly converges to a $\log$-$\Gamma$ distribution when $L\rightarrow\infty$.
@@
In general, the coefficients $A_{L}$ are expected to grow factorially [38] and, hence, the perturbative series, such as $\mathcal{A}(\hbar)=\sum_{L\geq 0}A_{L}\hbar^{L}$, to be divergent if $\hbar\neq 0$. To still utilize the information of the coefficients $A_{L}$ and to rigorously compute a quantity such as $\mathcal{A}(\hbar)$ for nonzero $\hbar$, resummation techniques are needed. As input data, these techniques need information about the $L\rightarrow\infty$ asymptotic behavior of $A_{L}$. For this reason, this asymptotic behavior is of high conceptual and practical value for the perturbative quantum field theory framework [39, 40, 41, 42, 43]. One concrete application is the prediction of critical exponents of various interesting universality classes via the Wilson–Fisher approach [2] (see also [44] for a recent discussion of the relevance of large-order contributions in this domain). In §3, we tabulate our 17 loop order results for the value of the primitive contribution to the $\phi^{4}$ beta function. We then extrapolate these values to the limit $L\rightarrow\infty$ and discuss the relation to classic conjectures on the asymptotic growth rate of the $\phi^{4}$ beta function in the MS scheme (e.g. to [45, 22, 24, 25]). We conclude in §4.
@@ 2.1 The primitive contribution to the $\phi^{4}$ beta function
In four dimensions, the superficial degree of divergence of a scalar Feynman graph is given by $\omega(G)=|E_{G}|-2L(G)$, where $|E_{G}|$ is the number of edges of the graph and $L(G)$ its loop number [46]. A graph $G$ is primitive divergent in $\phi^{4}$ theory if it is 1PI, and $\omega(\gamma)>0$ for each proper subgraph $\gamma\subsetneq G$, while $\omega(G)=0$. The last conditions ensure the graph has an overall logarithmic divergence and no subdivergences in four-dimensional spacetime. As usual, we will consider the legs of the graphs fixed or equivalently distinguishable. The condition $\omega(G)=0$ in $\phi^{4}$ theory in four dimensions implies that the graph has precisely four external legs.
@@
In this article, we focus on the primitive contribution to the $\phi^{4}$ beta function (see [23, Appendix B], whose notation we follow, for details). This contribution is renormalization scheme independent. We may express this contribution in terms of period Feynman integrals,111 The convential shift by one in $\beta^{\mathrm{prim}}_{L+1}$ is due to the fact that $L$-loop vertex diagrams contribute to the $g^{L+1}$ coefficient of the beta function in the coupling $g$.
@@ 2.2 A probabilistic approach to the sum over graphs
For our purposes, it is convenient to sample graphs using this non-uniform distribution. One reason is that it fits naturally with the perturbative expansion in Eq. (1), where each graph is weighted by $1/|\operatorname{Aut}(G)|$. Further, it is relatively straightforward to implement a sampling algorithm for this distribution, as we will show below. Moreover, we avoid the cumbersome explicit computation of the cardinality of the automorphism group $\operatorname{Aut}(G)$ (see, e.g., [48]) entirely.
@@
As the legs are distinguishable (which can be realized on the computer by numbering them, for example), the graph returned by the algorithm will have four distinguishable legs. It follows from a simple combinatorial argument and the orbit stabilizer theorem (see, e.g., Lemma 1 of [49, §5.1]) that the algorithm above randomly produces graphs $G$ with probability $p(G)$ from (6). Combining (5) and (6), gives
@@ 3.1 Histograms of Feynman integrals at large loop order
Each integral was evaluated using the non-parallelized version of the tropical sampling implementation published with [26]. At 17 loops a single integration required about 128 GB of RAM and an average runtime of $1.6\cdot 10^{4}$ seconds. Roughly two-thirds of this time was spent on the pre-processing of the integral, i.e., constructing the tropical measure (see [26, Sec. 7.2] for a precise definition), while the remaining third is spent on the Monte Carlo sampling. At lower loop order the pre-processing time and memory requirements decrease substantially. The ancillary material includes, for each integral, the time spent on both the pre-processing and the sampling steps.
@@
where we sum over all period samples $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}$ at a specific loop order. This likelihood function is readily derived by multiplying $N$ copies of the function that multiplies the measure in (2) and taking the logarithm. As the number of samples $N$ is large, we can estimate the uncertainties of these parameters by approximating the likelihood function near its maximum with a Gaussian. The resulting parameters with uncertainties are listed in Table 2. The uncertainties of the fit parameters were extremely large for $L\in\{8,9\}$. So, we discarded these fits. The fitted distributions are depicted as red curves in Figure 1 and Figure 2.
We were not able to give a clear conjecture for the concrete dependence of the parameters $\alpha$, $\lambda$, and $\mathcal{P}_{0}$ on the loop order $L$. Finding a formula for this explicit dependence, at least in the limit $L\rightarrow\infty$, would be highly beneficial. The dependence is illustrated in Figure 3. To give a qualitative assessement, we fitted the parameters $\alpha$, $\lambda$, and $\log\mathcal{P}_{0}$ using the shifted power-law function:
$f(L;a,b,c)=aL^{b}+c.$ (8)
The resulting fit parameters $a$, $b$, and $c$ are summarized in Table 1 , and the corresponding fits are shown as solid curves in Figure 3. Unfortunately, the fits do not provide a clear and sufficiently convincing picture for the dependence in the $L\rightarrow\infty$ limit.
$a$ $b$ $c$
$\alpha(L)=f(L;a,b,c)$ $1.0(1.8)\cdot 10^{12}$ 10.92(71)-10.92(71) $6.586(77)$
$\lambda(L)=f(L;a,b,c)$ $1.0(1.2)\cdot 10^{8}$ 7.21(51)-7.21(51) $3.849(74)$
$\log\mathcal{P}_{0}(L)=f(L;a,b,c)$ $80(67)$ $0.084(58)$ 90(68)-90(68)
Table 1: Best-fit parameters $a$, $b$, and $c$ for the $\log$-$\Gamma$ distribution parameters as shifted power-law functions (8) in the loop order $L$.
@@
We checked Conjecture 1 quantitatively using Pearson’s $\chi^{2}$ test: Let $O_{i}$ be the number of evaluated period Feynman integrals that fall into the $i$-th percentile of the distribution (2) with the fitted maximum likelihood parameters at the respective loop order. The expectation value of the random variable $O_{i}$ is $N/100$. So, under the hypothesis that our data follows (2), the quantity $\chi^{2}=\frac{100}{N}\sum_{i}(O_{i}-\frac{N}{100})^{2}$ is expected to follow a $\chi^{2}$-distribution with mean 1003=97100-3=97, as three parameters are fitted. Table 2 shows that the ratio $\chi^{2}/97$ approaches $1$ with increasing loop order, consistent with Conjecture 1. The trend exhibited in Table 2 is further supported by the computed $p$-value at $L=17$, defined as the probability of obtaining a value of the test statistic $\chi^{2}$ at least as large as the one observed, assuming the fitted $\log$-$\Gamma$ distribution accurately describes the data. At $L=17$, we find $p=0.087$, indicative of a statistically non-significant deviation from the expected distribution. The observed convergence of the ratio $\chi^{2}/97$ towards unity further reinforces that deviations diminish progressively with increasing loop order, strongly supporting Conjecture 1. Figure 2 illustrates how the distribution is approached with increasing loop number, providing further evidence for Conjecture 1.
@@
$L$ $\alpha$ $\lambda$ $\mathcal{P}_{0}$ $\frac{\chi^{2}}{97}$ $\langle\mathcal{P}(G)\rangle_{L}^{\rm CLT}$ $\langle\mathcal{P}(G)\rangle_{L}^{\log\text{-}\Gamma}$ $\beta_{L+1}^{\rm prim}$ NN 88 $7.449(7)\cdot 10^{2}$ $6.064(5)\cdot 10^{6}$ $99900$ 99 $1.644(2)\cdot 10^{3}$ $1.046(1)\cdot 10^{8}$ $99800$ $10$ $19.72(84)$ $10.6(2)$ $4.97(19)\cdot 10^{2}$ $273.7$ $3.504(5)\cdot 10^{3}$ $3.522(5)\cdot 10^{3}$ $1.890(3)\cdot 10^{9}$ $99900$ $11$ $10.69(27)$ $6.9(1)$ $1.36(3)\cdot 10^{3}$ $67.7$ $7.237(12)\cdot 10^{3}$ $7.303(14)\cdot 10^{3}$ $3.558(6)\cdot 10^{10}$ $99700$ $12$ $8.36(17)$ $5.53(6)$ $2.78(4)\cdot 10^{3}$ $19.3$ $1.460(3)\cdot 10^{4}$ $1.476(3)\cdot 10^{4}$ $6.998(14)\cdot 10^{11}$ $100000$ $13$ $7.38(13)$ $4.85(5)$ $5.28(6)\cdot 10^{3}$ $7.3$ $2.869(6)\cdot 10^{4}$ $2.900(7)\cdot 10^{4}$ $1.429(3)\cdot 10^{13}$ $98712$ $14$ $6.88(11)$ $4.42(4)$ $9.58(10)\cdot 10^{3}$ $4.6$ $531(14)\cdot 10^{4} 5.5$$588(15)\cdot 10^{4}$ $3.036(7)\cdot 10^{14}$ $100000$ $15$ $6.83(10)$ $4.24(4)$ $1.67(2)\cdot 10^{4}$ $2.5$ $1.039(3)\cdot 10^{5}$ $1.046(3)\cdot 10^{5}$ $6.654(18)\cdot 10^{15}$ $100000$ $16$ $6.688(96)$ $4.08(4)$ $92(3)\cdot 10^{4} 2.2$2.2 $1.910(6)\cdot 10^{5}$ $1.916(6)\cdot 10^{5}$ $1.504(4)\cdot 10^{17}$ $99650$ $17$ $6.41(12)$ $3.87(5)$ $5.14(7)\cdot 10^{4}$ $1.2$ $3.489(17)\cdot 10^{5}$ $3.495(17)\cdot 10^{5}$ $3.556(17)\cdot 10^{18}$ $44027$
@@
Table 2: Fit parameters $\alpha,\lambda,$ and $\mathcal{P}_{0}$ for the period Feynman integral distribution in $\phi^{4}$ theory, the normalized $\chi^{2}$ value resulting from Pearson’s $\chi^{2}$ test, the estimated average value of the period Feynman integral $\langle{\mathcal{P}}(G)\rangle_{L}^{\rm CLT}$ using the central limit theorem, $\langle{\mathcal{P}}(G)\rangle_{L}^{\log\text{-}\Gamma}$ using the $\log$-$\Gamma$ parameters and Eq. (9), the estimated value of the primitive beta function coefficient, and the number $N$ of Feynman integrals evaluated at each loop order.
Figure 3: Fitted parameters $\alpha$, $\lambda$, and $\log(\mathcal{P}_{0})$ of the $\log$-$\Gamma$ distribution (2) as a function of the loop order $L$. The purple curves are fitted according to the shifted power-law (8).